Solve the equation 25x2 - 64 = 0 using the quadratic formula. (Notice b = 0. Enter solutions from smallest to largest.)
x = ?
x = ?
Please show me the steps to get the steps.

Answer by mananth(12270)

(5x)^2-(8)^2=0
(5x+8)(5x-8)=0
5x=-8
x=-8/5
OR
5x-8=0
5x=8
x=8/5